# Estimating non-linear functionals of trawl processes

### v2

@@ Abstract
Trawl processes is a family of continuous-time, infinitely divisible, stationary processes whose correlation structure is entirely characterized by its so-called trawl function. This paper investigates the problem of estimating non-linear functionals of a trawl function under an in-fill and a long-span sampling scheme. Specifically, building on the work of [17], we introduce non-parametric estimators for functionals of the type $\Psi_{t}(g)=\int_{0}^{t}g(a(s))\mathrm{d}s$ and $\Lambda_{t}(g)=\int_{t}^{\infty}g(a(s))\mathrm{d}s$ where $a$ represents the trawl function of interest and $g$ a non-linear test function. We show that our estimator for $\Psi_{t}(g)$ is consistent and asymptotically Gaussian regardless of the memory of the process. We further demonstrates that the same phenomenon occurs for the estimation of $\Lambda_{t}(g)$ as long as $g(x)=\mathrm{O}(\lvert x\rvert^{p})$, as $x\to 0$, for some $p>3$. Additionally, we illustrate how our results can be used to test the presence of $T$-dependent data that is robust to persistence.
1 Introduction
Trawl processes form a subclass of continuous-time, strictly stationary, and infinitely divisible processes. Specifically, a trawl process $X=(X_{t})_{t\in\mathbb{R}}$ is constructed by evaluating a Lévy basis – also known as infinitely divisible independently scattered random measure – $L$ over a set $A_{t}$, known as trawl set, which is typically a subset of $\mathbb{R}^{2}$. A trawl process, by its design, offers highly flexible autocorrelation structure and the ability to produce a wide range of marginal distributions, all within the framework of infinitely divisible distributions. In fact, the distribution of $X_{t}$ is completely determined by $L$, while the autocorrelation function of $X$ is fully described by the so-called trawl function $a$. This relationship is often expressed through the identity
@@ 1 Introduction
These features make trawl processes suitable for modeling data with stylized facts such as non-Gaussianity, heavy tails, jumps, and persistence. This is reflected by their increasingly popularity for modeling complex temporal phenomena across disciplines such as finance ([3, 4, 20, 23, 24]) and physics ([9, 14, 18]).
Regarding statistical inference for trawl processes, various parametric approaches have been explored in the literature. See for instance [1, 3, 4, 19]. In this context, the work in [17] represents the first to address non-parametric estimation of the trawl function and serves as the starting point for the present study. Specifically, the authors in that paper showed that, given $n$ equidistant observations of $X$, say $(X_{i\Delta_{n}})_{i=0}^{n}$, the trawl function $a$ can be consistently estimated by
@@
for some non-negative constant $\mathfrak{K}_{4}$. In the same work, the authors proposed to use $\check{a}$ once again along with a “Riemann sum” approach to estimate the quadratic functionals of $a$ appearing in the definition of $\sigma_{a}^{2}(t)$, e.g. $\int_{0}^{\infty}a(s)^{2}\mathrm{d}s$. This procedure has some limitations. First, it relies on a tuning parameter whose choice and finite-sample performance are unclear. Second, the estimation error of $\int_{0}^{\infty}a(s)^{2}\mathrm{d}s$ is dependent of the behaviour of $\check{a}(t)$ for $t$ large, which could result on a statistic that is very sensitive to the memory of the underlying stochastic process. In this paper, we address these issues by considering the general problem of estimating functionals of the type
$\Psi_{t}(g)=\int_{0}^{t}g(a(s))\mathrm{d}s;\,\,\Lambda_{t}(g):=\int_{t}^{\infty}g(a(s))\mathrm{d}s,$
@@
using a double-asymptotic scheme based on both long-span and high-frequency observations of the underlying trawl process $X$. The function $g$ is assumed to be continuous and of polynomial growth.
Under minimal integrability conditions on $X$, we provide a functional law of large numbers for $\Psi^{n}(g)$. With stronger assumptions on the trawl function, we establish a functional Central Limit Theorem (CLT from now on) at the rate $\sqrt{n\Delta_{n}}$, applicable to test functions that are twice continuously differentiable. Additionally, in line with [17], we show that this convergence is independent of the memory of $X$. In contrast, the asymptotic behaviour of $\Lambda^{n}(g)$ is more subtle. First, we demonstrate that, in general, $\Lambda_{t}^{n}(g)$ is not consistent for $\int_{t}^{\infty}g(a(s))\mathrm{d}s$ unless $g(x)=\mathrm{O}(\lvert x\rvert^{p})$, as $x\to 0$, for some $p>2$. In the latter situation with $p>3$, $\Lambda_{t}^{n}(g)$ also fulfils a CLT analogous to that of $\Psi^{n}(g)$. Secondly, to mitigate the inconsistency of $\Lambda^{n}(g)$ we alter its definition (see equation (13)) following the approach suggested in [17]. The resulting estimator $\bar{\Lambda}^{n}(g)$ is consistent and satisfies a CLT at rate $\sqrt{n\Delta_{n}}$. However, this modification requires a tunning parameter and its second-order asymptotics are now dependent on the memory of the process and the sample scheme. Finally, to illustrate the usefulness of our results, we propose a method for testing the presence of $T$-dependent data that is robust to long memory.
@@
The remainder of the paper is organized as follows. In Section 2 we introduce the main mathematical concepts and and notation that will be used through the paper. Furthermore, the definitions of trawl processes and its basic properties are also included in this part. In Section 3 we formulate the main results of the paper, concerning the asymptotics of the functionals $\Psi^{n}(g)$ and $\Lambda^{n}(g)$, along with our new test for $T$-dependence. Due to the technical nature of our proofs, most of them are deferred to Section 4.
2 Preliminaries
This part focus on introducing the basic notation and concepts that will be used during this paper. Throughout the following sections $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ will denote a complete probability space. As usual, a function $f:\mathbb{R}\rightarrow\mathbb{R}$ is said to be of class $\mathcal{C}^{N}$, $N\in\mathbb{N}$, if it is $N$-times continuously differentiable. In this case, $f^{(k)}$ will denote the $k$th derivative of $f$ with the convention that $f^{(0)}=f$. Often, we will write $f^{\prime}$ or $\dot{f}$ instead of $f^{(1)}$. The symetrization of a function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is defined and denoted as $\tilde{f}(x,y):=f(x,y)+f(y,x)$.
@@ 2 Preliminaries
We recall to the reader that a real-valued random field $L=\{L\left(A\right):A\in\mathcal{B}_{b}(\mathbb{R}^{d})\}$ on $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ is called a homogeneous Lévy basis if it is an infinitely divisible (ID for short) independently scatterred random measure such that
$\mathbb{E}(\exp(\mathbf{i}zL(A)))=\exp(\mathrm{Leb}(A)\psi(z)),\,\,\forall\,A\in\mathcal{B}_{b}(\mathbb{R}^{d}),z\in\mathbb{R},$ (2)
where $\mathrm{Leb}$ and $\mathcal{B}_{b}(\mathbb{R}^{d})$ denote the Lebesgue measure on $\mathbb{R}^{d}$ and the bounded Borel sets of $\mathbb{R}^{d}$, respectively. Furthermore,
@@
with $\gamma\in\mathbb{R},$ $b\geq 0$ and $\nu$ a Lévy measure, i.e. $\{$$\int_{\mathbb{R}\backslash\{0\}}(1\land\rvert x\rvert^{2})\nu(\mathrm{d}x)<\infty$ and $\nu$ charges 0 at 0{0}. We will refer to the ID random variable with characteristic triplet $\left(\gamma,b,\nu\right)$ as the Lévy seed of $L$, and it will be denoted by $L^{\prime}$. As it is costumary, $\left(\gamma,b,\nu\right)$ will be called the characteristic triplet of $L$ and $\psi$ its characteristic exponent. We will say that $L$ has unit variance if $\mathrm{Var}(L^{\prime})=1$.
@@
where $A_{t}:=A+(t,0)$, is called trawl process. To avoid trivial situations, we will always assume that $\mathrm{Leb}(A)>0$ or equivalently $a$ cannot be equal to 0 almost everywhere. From now on, we will allude to $A$ and $a$ as trawl set and trawl function, respectively. Furthermore, $L$ will termed as as the background driving Lévy basis.
@@
3 Estimating non-linear functionals of trawl functions
In what is left of this work $X$ will denote a trawl process with trawl function $a$ and whose background driving Lévy basis has unit variance. As already discussed in the introduction, our main goal is to estimate the functionals
@@ 3 Estimating non-linear functionals of trawl functions
when equidistant observations of $X$, say $(X_{i\Delta_{n}})_{i=0}^{n}$, are available. The sample scheme considered in this work is in-fill and long-span, i.e.
@@
To do so, we modify slightly the definition of $\check{a}$ in (1) as follows:
@@
for some $l=0,\ldots,n-1$ and where $\delta_{k}X:=X_{(k+1)\Delta_{n}}-X_{k\Delta_{n}}$. It is not difficult to see that $\hat{a}(t)$ and $\check{a}(t)$ are asymptotically equivalents. The estimators for the functionals appearing in (7) are constructed in a natural way as
@@
$\Lambda_{t}^{n}(g)$= $\Delta_{n}\sum_{l=[t/\Delta_{n}]}^{n-1}g(\hat{a}(l\Delta_{n}));0\leq t\leq(n-1)\Delta_{n}.$
3.1 Limit theorems for $\Psi^{n}(g)$ and $\Lambda^{n}(g)$
In this part we describe first and second order asymptotics for the estimators introduced in (10). We start discussing consistency of such statistics. We recall the reader that a real-valued function $f:\mathbb{R}\rightarrow\mathbb{R}$ is said to be of polynomial growth of order $q\geq 0$ if there is some $C>0$, such that
@@ 3.1 Limit theorems for $\Psi^{n}(g)$ and $\Lambda^{n}(g)$
The reader should keep in mind that our sampling scheme satisfies (8) and this will be always assumed to hold for the rest of this paper. Under this set-up we have the following consistency result for $\Psi^{n}(g)$.
@@
Let us now turn our attention to $\Lambda^{n}(g)$. To guarantee integrability of the mapping $s\mapsto g(a(s))$, ergo the well-definiteness of $\Lambda_{t}(g)$, we need to restrict the type of test functions to be considered. For a given $p,q\geq 0$, $d\in\mathbb{N}$, we will write $\mathfrak{C}_{p,q}^{d}$ for the family of functions of class $\mathcal{C}^{d}$ satisfying that $g^{(j)}(0)=0$ for $j=0,1,\ldots,d-1$ and that $g^{(d)}(x)=\mathrm{O}(\rvert x\rvert^{p})$ as $x\rightarrow 0$ and $g^{(d)}(x)=\mathrm{O}(\rvert x\rvert^{q})$ as $x\rightarrow+\infty$. Obviously the power function $g(x)=\rvert x\rvert^{\alpha}$, for $\alpha>1$ belongs to $\mathfrak{C}_{p,q}^{[\alpha]}$ with $p=q=\alpha-[\alpha]$.
@@
Note that if $g\in\mathfrak{C}_{p,q}^{d}$, an induction argument along with the Mean-Value Theorem imply that $g^{(j)}(x)=\mathrm{O}(\rvert x\rvert^{d-j+p})$, as $x\rightarrow 0$, for all $j=1,\ldots,d$. Thus, in view that $a$ is decreasing and bounded by $a(0),$ we can find a constant $C=C(a(0))>0$, such that for all $s\geq 0$
$\max_{j=0,1,\ldots,d-1}\rvert g^{(j)}(a(s))\rvert\leq Ca(s)^{d-j+p},\,\,\,j=1,\ldots,d.$ (11)
@@
Let $g\in\mathfrak{C}_{p,q}^{1}$ with $p,q>1$ and assume that $\mathbb{E}(\rvert L^{\prime}\rvert^{2(q\lor p+1)})<\infty$. Then, as $n\rightarrow\infty$, $\Lambda^{n}(g)\overset{u.c.p}{\rightarrow}\Lambda(g).$
The previous result does not cover the quadratic case, i.e. when $p\land q=1$. This is because in this situation $\Lambda^{n}(g)$ might be biased as the following result shows.
@@
To deal with the previous situation, we proceed as in [17] and modify $\Lambda^{n}(g)$ as follows: We introduce a sequence $N_{n}\in\mathbb{N}$, $N_{n}\leq n$, such that
@@
Consider $N_{n}\in\mathbb{N}$ as in (12). Let $p,q\geq 1$ and assume that $\mathbb{E}(\rvert L^{\prime}\rvert^{2(q\lor p+1)})<\infty$. Then for all $g\in\mathfrak{C}_{p,q}^{1}$, $\bar{\Lambda}_{t}^{n}(g)\overset{u.c.p}{\rightarrow}\Lambda(g)$, as $n\rightarrow\infty.$
We now proceed to discuss second order asymptotics for our statistics. As usual, to achieve this, stronger assumptions need to be imposed. We follow essentially the set-up of [17].
@@
Before presenting the Central Limit Theorems for $\Psi^{n},\Lambda^{n}$, and $\bar{\Lambda}^{n}$ let us introduce some functions that will serve for describing their asymptotic variance. For $s,r\geq 0$ and $\mathfrak{K}_{4}:=\int x^{4}\nu(dx)$, we put
@@
Let Assumption 1 hold. If $g:\mathbb{R}\rightarrow\mathbb{R}$ is of class $\mathcal{C}^{2}$, with $g^{(2)}$ being of polynomial growth of order $q\geq 0$ and $\mathbb{E}(\rvert L^{\prime}\rvert^{2(q+2)\lor r_{0}})<\infty$ for some $r_{0}>4$, then as $n\rightarrow\infty$
@@
where $(Z_{t})_{t\geq 0}$ is a centered Gaussian process with
$\mathbb{E}(Z_{t}Z_{s})=\int_{0}^{t}\int_{0}^{s}\dot{g}(a(u))\Sigma_{a}(u,r)\dot{g}(a(r))\mathrm{d}r\mathrm{d}u,\,\,\,t,s\geq 0,$
@@
$\Sigma_{a}(u,r)$ $=\sigma_{1}(u,r)+\tilde{\sigma}_{2}(u,r)+\tilde{\sigma}_{3}(u,r).$ (14)
The second order asymptotics for the functionals $\Lambda^{n}$ and $\bar{\Lambda}^{n}$ will be much alike to the one presented in the previous theorem. However, we will need to impose certain restrictions on the window $N_{n}$ that will help us to control the discretization error associated to $\bar{\Lambda}^{n}$ as well as the memory of $X$. Specifically, we will assume that:
@@
Let Assumption 1 hold. Suppose that $g\in\mathfrak{C}_{p,q}^{2}$ and that for some $r_{0}>4$ we have that $\mathbb{E}(\rvert L^{\prime}\rvert^{2(q\lor p+2)\lor r_{0}})<\infty$. Then the following holds:
@@
    where $(Z_{t})_{t\geq 0}$ is a centered Gaussian process with
    $\mathbb{E}(Z_{t}Z_{s})=\int_{t}^{\infty}\int_{s}^{\infty}\dot{g}(a(u))\Sigma_{a}(u,r)\dot{g}(a(r))\mathrm{d}r\mathrm{d}u,$
@@
The following remarks are in order
@@
    We are able to show that in the limiting case $p\wedge q=1$ – recall that when this is not the case, in general, $\Lambda^{n}(g)$ is not consistent – the sequence $\sqrt{n\Delta_{n}}(\Lambda^{n}(g)-\Lambda(g))$ is tight on $\mathcal{D}([0,T])$ but we were not able to identify its limit. However, we believe that, just as in the case $g(x)=x^{2}$, the limit process is Gaussian but no longer centred.

### v3

@@ Abstract
Trawl processes are a family of continuous-time, infinitely divisible, stationary processes whose correlation structure is entirely characterized by their so-called trawl function. This paper investigates the problem of estimating non-linear functionals of a trawl function under in-fill and long-span sampling schemes. Specifically, building on the work of [19], we introduce non-parametric estimators for functionals of the type $\Psi_{t}(g)=\int_{0}^{t}g(a(s))\mathrm{d}s$ and $\Lambda_{t}(g)=\int_{t}^{\infty}g(a(s))\mathrm{d}s$, where $a$ represents the trawl function of interest and $g$ a non-linear test function. We show that our estimator for $\Psi_{t}(g)$ is consistent and asymptotically Gaussian regardless of the memory of the process. We further demonstrate that the same phenomenon occurs for the estimation of $\Lambda_{t}(g)$ as long as $g(x)=\mathrm{O}(\lvert x\rvert^{p})$, as $x\to 0$, for some $p>3$. Additionally, we illustrate how our results can be used to construct a test statistic robust to memory effects for the presence of $T$-dependent.
Keywords: Functional limit theorems; Infinitely divisible processes; Nonparametric estimation; Trawl processes.
1.  Introduction
Trawl processes form a subclass of continuous-time, strictly stationary, and infinitely divisible processes. Specifically, a trawl process $X=(X_{t})_{t\in\mathbb{R}}$ is constructed by evaluating a Lévy basis – also known as an infinitely divisible independently scattered random measure – $L$ over a set $A_{t}$, known as a trawl set, which is typically a subset of $\mathbb{R}^{2}$. The distribution of $X_{t}$ is completely determined by $L$, while the autocorrelation function of $X$ is fully described by the so-called trawl function $a$. This relationship is expressed through the identity
@@ 1.  Introduction
By design, a trawl process offers a highly flexible autocorrelation structure and produces a wide range of marginal distributions within the class of infinitely divisible distributions. These features make trawl processes suitable for modeling data with stylized facts such as non-Gaussianity, heavy tails, jumps, and persistence. This is reflected in their increasing popularity for modeling complex temporal phenomena across disciplines such as finance ([2, 4, 22, 25, 26]) and physics ([10, 16, 20]).
Regarding statistical inference for trawl processes, various parametric approaches have been explored in the literature. See for instance [1, 2, 4, 21]. In this context, the work in [19] is the first to address non-parametric estimation of the trawl function and serves as the starting point for the present study. Specifically, that paper proves that, given $n$ equidistant observations of $X$, say $(X_{i\Delta_{n}})_{i=0}^{n}$, the trawl function $a$ can be consistently estimated by
@@
for some non-negative constant $\mathfrak{K}_{4}$. In the same work, we proposed to use $\check{a}$ once again along with a “Riemann sum” approach to estimate the quadratic functionals of $a$ appearing in the definition of $\sigma_{a}^{2}(t)$, e.g. $\int_{0}^{\infty}a(s)^{2}\mathrm{d}s$. This procedure has some limitations. First, it relies on a tuning parameter whose choice and finite-sample performance are unclear. Second, the estimation error of $\int_{0}^{\infty}a(s)^{2}\mathrm{d}s$ is dependent on the behaviour of $\check{a}(t)$ for large $t$, which could result on a statistic that is very sensitive to the memory of the underlying stochastic process. In this paper, we address these issues by estimating functionals of the type
$\Psi_{t}(g)=\int_{0}^{t}g(a(s))\mathrm{d}s;\,\,\Lambda(g):=\int_{t}^{\infty}g(a(s))\mathrm{d}s,$
@@
using a double-asymptotic scheme based on both long-span and high-frequency observations of the underlying trawl process $X$. The function $g$ is assumed to be continuous.
Under minimal integrability conditions on $X$ and a growth condition on $g$, we provide a functional law of large numbers for $\Psi^{n}(g)$. With stronger assumptions on the trawl function, we establish a functional Central Limit Theorem (CLT hereafter)) at the rate $\sqrt{n\Delta_{n}}$, applicable to test functions that are twice continuously differentiable and of polynomial growth. Additionally, in line with [19], we show that this convergence is independent of the memory of $X$. In contrast, the asymptotic behaviour of $\Lambda^{n}(g)$ is more subtle. First, we demonstrate that, in general, $\Lambda_{t}^{n}(g)$ is not consistent for $\int_{t}^{\infty}g(a(s))\mathrm{d}s$ unless $g(x)=\mathrm{O}(\lvert x\rvert^{p})$, as $x\to 0$, for some $p>2$. In the latter situation with $p>3$, $\Lambda_{t}^{n}(g)$ also fulfils a CLT analogous to that of $\Psi^{n}(g)$. Secondly, to mitigate the inconsistency of $\Lambda^{n}(g)$ we alter its definition (see equation (13)) following the approach suggested in [19]. The resulting estimator $\bar{\Lambda}^{n}(g)$ is consistent and satisfies a CLT at rate $\sqrt{n\Delta_{n}}$. However, this modification requires a tuning parameter and its second-order asymptotics are dependent on the memory of the process and the sample scheme. Finally, to illustrate the usefulness of our results, we propose a method for testing the presence of $T$-dependent data using a statistic that is robust to long memory.
@@
The remainder of the paper is organised as follows. Section 2 introduces the main mathematical concepts and notation used throughout the paper and recalls the definition and basic properties of trawl processes. Section 3 presents the main results concerning the asymptotics of the functionals $\Psi^{n}(g)$ and $\Lambda^{n}(g)$, along with our new test for $T$-dependence. Due to the technical nature of the arguments, most proofs are deferred to Section 4. The paper also contains an appendix where we carry out the computations required to approximate the asymptotic variance appearing in our central limit theorems. The arguments done there are primarily algebraic and technical in nature, and the appendix may therefore be skipped on a first reading.
2.  Preliminaries
This part introduces the basic notation and concepts that will be used in this paper. Throughout the following sections $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ denotes a complete probability space. As usual, a function $f:\mathbb{R}\rightarrow\mathbb{R}$ is said to be of class $\mathcal{C}^{N}$, $N\in\mathbb{N}$, if it is $N$-times continuously differentiable. In this case, $f^{(k)}$ will denote the $k$th derivative of $f$, with the convention that $f^{(0)}=f$. Often, we will write $f^{\prime}$ instead of $f^{(1)}$. The symmetrization of a function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is defined and denoted as $\tilde{f}(x,y):=f(x,y)+f(y,x)$.
@@ 2.  Preliminaries
We recall that a real-valued random field $L=\{L\left(A\right):A\in\mathcal{B}_{b}(\mathbb{R}^{d})\}$ on $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ is called a homogeneous Lévy basis if it is an infinitely divisible (ID for short) independently scattered random measure such that
$\mathbb{E}\big(\exp(\mathbf{i}zL(A))\big)=\exp\big(\mathrm{Leb}(A)\psi(z)\big),\,\,\forall\,A\in\mathcal{B}_{b}(\mathbb{R}^{d}),z\in\mathbb{R},$ (2)
where $\mathrm{Leb}$ and $\mathcal{B}_{b}(\mathbb{R}^{d})$ denote the Lebesgue measure on $\mathbb{R}^{d}$ and the Borel sets of $\mathbb{R}^{d}$ with finite Lebesgue measure, respectively. Furthermore,
@@
with $\gamma\in\mathbb{R},$ $b\geq 0$ and $\nu$ a Lévy measure, i.e. $\{$$\int_{\mathbb{R}\backslash\{0\}}(1\land\rvert x\rvert^{2})\nu(\mathrm{d}x)<\infty$ and $\nu$ does not charge 0{0}. We will refer to the ID random variable with characteristic triplet $\left(\gamma,b,\nu\right)$ as the Lévy seed of $L$, and it will be denoted by $L^{\prime}$. As it is customary, $\left(\gamma,b,\nu\right)$ will be called the characteristic triplet of $L$ and $\psi$ its characteristic exponent. We will say that $L$ has unit variance if $\mathrm{Var}(L^{\prime})=1$.
@@
where $A_{t}:=A+(t,0)$, is called a trawl process. To avoid trivial situations, we will always assume that $\mathrm{Leb}(A)>0$, or equivalently that $a$ is not equal to 0 almost everywhere. From now on, we will allude to $A$ and $a$ as trawl set and trawl function, respectively. Furthermore, $L$ is termed the background driving Lévy basis.
@@
3.  Estimating non-linear functionals of trawl functions
In what is left of this work, $X$ will denote a trawl process with trawl function $a$ and whose background driving Lévy basis has unit variance. As discussed in the introduction, our main goal is to estimate the functionals
@@ 3.  Estimating non-linear functionals of trawl functions
when equidistant observations of $X$, say $(X_{i\Delta_{n}})_{i=0}^{n}$, are available. The sample scheme considered in this work is an in-fill and long-span, i.e.
@@
To do so, we slightly modify the definition of $\check{a}$ in (1) as follows:
@@
for some $l=0,\ldots,n-1$ and where $\delta_{k}X:=X_{(k+1)\Delta_{n}}-X_{k\Delta_{n}}$. It is not difficult to see that $\hat{a}(t)$ and $\check{a}(t)$ are asymptotically equivalent. The estimators for the functionals appearing in (7) are constructed as
@@
$\Lambda_{t}^{n}(g)$= $\Delta_{n}\sum_{l=[t/\Delta_{n}]}^{n-1}g(\hat{a}(l\Delta_{n}));0\leq t\leq(n-1)\Delta_{n},$
where $[x]$ denotes the integer part of $x\in\mathbb{R}$.
3.1. Limit theorems for $\Psi^{n}(g)$ and $\Lambda^{n}(g)$
In this part we describe first and second order asymptotics for the estimators introduced in (10). We start by discussing the consistency of these statistics. We recall that a real-valued function $f:\mathbb{R}\rightarrow\mathbb{R}$ is said to be of polynomial growth of order $q\geq 0$ if there is some $C>0$, such that
@@ 3.1. Limit theorems for $\Psi^{n}(g)$ and $\Lambda^{n}(g)$
The reader should keep in mind that our sampling scheme satisfies (8) and this is assumed throughout the rest of the paper. Under this set-up we have the following consistency result for $\Psi^{n}(g)$.
@@
Let us now turn our attention to $\Lambda^{n}(g)$. To guarantee integrability of the mapping $s\mapsto g(a(s))$, and hence the well-definedness of $\Lambda_{t}(g)$, we need to restrict the type of test functions. For a given $p,q\geq 0$, $d\in\mathbb{N}$, we will write $\mathfrak{C}_{p,q}^{d}$ for the family of functions of class $\mathcal{C}^{d}$ satisfying that $g^{(j)}(0)=0$ for $j=0,1,\ldots,d-1$ and that $g^{(d)}(x)=\mathrm{O}(\rvert x\rvert^{p})$ as $x\rightarrow 0$ and $g^{(d)}(x)=\mathrm{O}(\rvert x\rvert^{q})$ as $x\rightarrow+\infty$. For example, the power function $g(x)=\rvert x\rvert^{\alpha}$, for $\alpha>1$ belongs to $\mathfrak{C}_{p,q}^{[\alpha]}$ with $p=q=\alpha-[\alpha]$.
@@
Note that if $g\in\mathfrak{C}_{p,q}^{d}$, an induction argument along with the Mean-Value Theorem imply that $g^{(j)}(x)=\mathrm{O}(\rvert x\rvert^{d-j+p})$, as $x\rightarrow 0$, for all $j=1,\ldots,d$. Thus, in view of the fact $a$ is decreasing and bounded by $a(0),$ we can find a constant $C=C(a(0))>0$, such that for all $s\geq 0$
$\rvert g^{(j)}(a(s))\rvert\leq Ca(s)^{d-j+p},\,\,\,j=1,\ldots,d.$ (11)
@@
Let $g\in\mathfrak{C}_{p,q}^{1}$ with $p,q>1$ and assume that $\mathbb{E}(\rvert L^{\prime}\rvert^{2((q\lor p)+1)})<\infty$. Then, as $n\rightarrow\infty$, $\Lambda^{n}(g)\overset{u.c.p}{\rightarrow}\Lambda(g).$
The previous result does not cover the quadratic case, i.e. when $p\land q=1$. This is because in this situation $\Lambda^{n}(g)$ might be biased, as the following result shows.
@@
To deal with the previous situation, we proceed as in [19] and modify $\Lambda^{n}(g)$ as follows: We introduce a tuning parameter $N_{n}\in\mathbb{N}$, $N_{n}\leq n$, such that
@@
Consider $N_{n}\in\mathbb{N}$ as in (12). Let $p,q\geq 1$ and assume that $\mathbb{E}(\rvert L^{\prime}\rvert^{2((q\lor p)+1)})<\infty$. Then for all $g\in\mathfrak{C}_{p,q}^{1}$, $\bar{\Lambda}_{t}^{n}(g)\overset{u.c.p}{\rightarrow}\Lambda(g)$, as $n\rightarrow\infty.$
We now discuss second-order asymptotics for our statistics. As usual, this requires stronger assumptions. We essentially follow the set-up of [19].
@@
Before presenting the Central Limit Theorems for $\Psi^{n},\Lambda^{n}$, and $\bar{\Lambda}^{n}$ let us introduce some functions that will serve to describe their asymptotic variance. For $s,r\geq 0$ and $\mathfrak{K}_{4}:=\int x^{4}\nu(dx)$, we put
@@
Let Assumption 1 hold. If $g:\mathbb{R}\rightarrow\mathbb{R}$ is of class $\mathcal{C}^{2}$ with $g^{(2)}$ being of polynomial growth of order $q\geq 0$, and $\mathbb{E}(\rvert L^{\prime}\rvert^{2(q+2)\lor r_{0}})<\infty$ for some $r_{0}>4$, then as $n\rightarrow\infty$,
@@
where $(Z_{t})_{t\geq 0}$ is a centred Gaussian process with
$\mathbb{E}(Z_{t}Z_{s})=\int_{0}^{t}\int_{0}^{s}g^{\prime}(a(u))\Sigma_{a}(u,r)g^{\prime}(a(r))\mathrm{d}r\mathrm{d}u,\,\,\,t,s\geq 0,$
@@
$\Sigma_{a}(u,r)$ $=\sigma_{1}(u,r)+\tilde{\sigma}_{2}(u,r)+\tilde{\sigma}_{3}(u,r),$ (14)
with $\tilde{\sigma}_{2}$ and $\tilde{\sigma}_{3}$ the symmetrizations (see Section 2) of $\sigma_{2}$ and $\sigma_{3}$, respectively.
The second order asymptotics for the functionals $\Lambda^{n}$ and $\bar{\Lambda}^{n}$ will be much alike to those presented in the previous theorem. However, we need to impose certain restrictions on the window $N_{n}$ to control the discretization error associated with $\bar{\Lambda}^{n}$, as well as the memory of $X$. Specifically, we assume that:
@@
Let Assumption 1 hold. Suppose that for some $r_{0}>4$, $\mathbb{E}(\rvert L^{\prime}\rvert^{2((q\lor p)+2)\lor r_{0}})<\infty$, and that $g\in\mathfrak{C}_{p,q}^{2}$. Then the following holds:
@@
    where $(Z_{t})_{t\geq 0}$ is a centred Gaussian process with
    $\mathbb{E}(Z_{t}Z_{s})=\int_{t}^{\infty}\int_{s}^{\infty}g^{\prime}(a(u))\Sigma_{a}(u,r)g^{\prime}(a(r))\mathrm{d}r\mathrm{d}u,$
@@
The following remarks are in order:
@@
    We are able to show that in the limiting case $p\wedge q=1$ – recall that when $p\wedge q<1$, in general, $\Lambda^{n}(g)$ is not consistent – the sequence $\sqrt{n\Delta_{n}}(\Lambda^{n}(g)-\Lambda(g))$ is tight on $\mathcal{D}([0,T])$ but we were not able to identify its limit. However, we believe that, just as in the case $g(x)=x^{2}$, the limit process is Gaussian but no longer centred.